OpenStudy (anonymous):

Your average (arithmetic mean) score on four exams is 82. If your professor drops the lowest exam score then your average will be 86. What was your lowest score?

OpenStudy (amistre64):

A+L ---- = 82 4 A/3 = 86 We can solve for A and then determine the value for the lowest score(L)

OpenStudy (amistre64):

id say L = 82(4) -86(3)

OpenStudy (anonymous):

Thanks,how did you derive that formula.

OpenStudy (amistre64):

i had to name the stuff to organize it; I figured A could be the rest of the grades; and L seemed a good candidate for the lowest; the rest of it is just utilizing what it means to take an average

OpenStudy (anonymous):

Good Answer!!!

OpenStudy (amistre64):

thnx :)